OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 2) = 12

mathslover (mathslover):

\[\huge{4x-4-x-2=12}\] \[\huge{3x-6=12}\] \[\huge{3x=18}\] \[\huge{x=\frac{18}{3}=6}\]

OpenStudy (anonymous):

Hm.....

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Helppp

mathslover (mathslover):

where r u having problem

mathslover (mathslover):

have a look at my tutorials